RSS Feed
News
Dec
20
DSL outage between 14.40 & 14.47 – 20/12/2010
Posted by Thomas Brown on 20 December 2010 05:36 PM

Background:

At 14.40 today we experienced a crash on the main Ethernet card within one of Broadband Remote Access Servers (BRAS).

Impact:

Customers with @n4dsl usernames lost connectivity between 14.40 to 14.47

Actions:
The crash info will be sent to TAC for analysis and we will take appropriate action based on their response.

Server Centre apologises for this issue and are doing everything possible to resolve this issue as quickly as we can.

Further information will be posted as it becomes available.


Read more »



Dec
20
Packet loss on @n4dsl connections
Posted by Thomas Brown on 20 December 2010 05:30 PM

Update 16.55 20/12/10: We have now received confirmation that the issue at the Faraday BT POP is now completely resolved. An Incident Report will be created over the course of the next couple of days , once we have all the facts for the incident. Contact your account manager or technical support should you wish to receive a copy of this report .

Update 16.45 20/12/10: We’ve received an update to say that the Faraday POP issue is improving and traffic is being restored. Therefore we expect to see improvements on the current situation.

Update 16.10 20/12/10: The BT issue at the Faraday POP is affecting a core piece of the network which BT engineers are working on at the moment in order to resolve. This is resulting in packet loss, slow speeds and intermittent connectivity for many DSL customers across the UK. This is a fairly significant issue for them to resolve and at present they are unable to confirm precisely when it will be fixed.

Update 13.45 20/12/10: The issue has been traced back to the BT Faraday POP in London. BT are currently working on the issue.

Update 13.45 20/12/10: The issue has been traced back to the BT Faraday POP in London. BT are currently working on the issue.

Update 13.45 20/12/10: The issue has been traced back to the BT Faraday POP in London. BT are currently working on the issue.

Update 13.45 20/12/10: The issue has been traced back to the BT Faraday POP in London. BT are currently working on the issue.

Update 13.45 20/12/10: The issue has been traced back to the BT Faraday POP in London. BT are currently working on the issue.

Update 13.45 20/12/10: The issue has been traced back to the BT Faraday POP in London. BT are currently working on the issue.

Background:

At approximately 12.20 today we received reports from customers stating that they were experiencing packet loss on their @n4dsl DSL connections. Engineers investigated and identified that this packet loss is affecting many customers connections. This issue has been escalated to our upstream provider.

Impact:

Poor performance / Packet loss on @n4dsl connections

Actions:
Escalated to upstream provider. They have confirmed that they have a core network related issue and have engineers investing.

Server Centre apologises for this issue and are doing everything possible to resolve this issue as quickly as we can.

Further information will be posted as it becomes available.

Regards,

Server Centre Technical Support


Read more »



Dec
11
Issues affecting DSL customers
Posted by Adam Heavens on 11 December 2010 04:14 PM

At approximately 2:20pm on the 11th December, our operations centre noticed a number of DSL connections dropping off the network. Senior engineers investigated these drops and the problem was found to be at KCom, one of our upstream providers.

We raised this fault with KCom who were already aware of an issue within their network.

We monitored the situation closely throughout the afternoon and could see intermittent connectivity from our IP Stream DSL connections.

At approximately 4:15pm most of the connections had re-established and stabilized.

We were informed by KCom at 5:13pm that they had restored connectivity to all connections. Further investigation by KCom’s senior engineers identified a hardware failure on a core device within their network.

Further information on this incident can be found on KCom’s website at http://status.mistral.net/status.php?appPage=Search&sStatus=*OPEN&appTask=View&appItem=1338&firstRow=0

Best regards

Server Centre Technical Support


Read more »